Data Trawler - Survey Details

Please login if you have access to particular applications.
Username:
Password:

Voyage details and related metadata, reports, events and data.

IN2019_E01    - is a Marine National Facility voyage

Title: Calibration, Trials and Upcoming Voyage User Equipment Testing

Period: 2019-07-19 22:30 to 2019-08-02 02:00 UTC

Duration: 13.15 days

From: Singapore To: Cairns

Region: Indian Ocean; Timor Sea; Arafura Sea; Coral Sea; Great Barrier Reef; Queensland coast northeast.

Ship: Investigator (RV) [details]

Description:

Project: Investigator - general vessel operations, gear trials, etc. [details] Primary


Voyage extents: 16° 50.6' S to 9° 07.0' S    132° 47.1' E to 146° 20.7' E Get voyage track shapefile Get CSV


Marlin Metadata:-

  • RV Investigator Radon Data - Selected Voyages (2014 onwards) [details]

  • RV Investigator Sea Carbon Dioxide (CO2) Data - Selected Voyages (2014 onwards) [details]

  • RV Investigator Voyage IN2019_E01 ADCP Data [details]

  • RV Investigator Voyage IN2019_E01 EK60 Echosounder [details]

  • RV Investigator Voyage IN2019_E01 EM122 Multibeam Echosounder Data [details]

  • RV Investigator Voyage IN2019_E01 EM710 MKII Multibeam Echosounder Data [details]

  • RV Investigator Voyage IN2019_E01 End of Voyage (EOV) Archive [details]

  • RV Investigator Voyage IN2019_E01 Gravity Data [details]

  • RV Investigator Voyage IN2019_E01 SBP120 Sub-bottom Profiler Data [details]

  • RV Investigator Voyage IN2019_E01 Sound Velocity Profile Data [details]

  • RV Investigator Voyage IN2019_E01 Underway UWY Data (Unprocessed) [details]

  • RV Investigator Voyage IN2019_E01 XBT Data [details]

  • Cruise Summary Report at SeaDataNet:-

  • View Cruise Summary Report at https://csr.seadatanet.org
  • Participants - on ship:- Click on column header to sort

    NameRoleOrganisation Project
    Jason Fazey Voyage Manager CSIRO MNF Voyage Management on RV Investigator

    This list includes those that may have only been on part of the voyage. See voyage reports on 'Reports and Maps' tab for details.

    MNF Reports

    Search for publications and reports from this and other voyages

    Data publishing

    This is a list of online datasets hosted by CSIRO. Other processed datasets hosted by collaborating organisations may be found on the 'Metadata' tab.
    Use the Get data link on the right to extract data using the NCMI Data Trawler.

    Data StreamData SummaryGet data
    ADCP Data
    Time (UTC): 2019-07-27 20:07 to 2019-08-02 11:08
    Latitude: -16.93 to -9.17
    Longitude: 132.97 to 146.26
    Ensembles: 1,763
    Get Data from Trawler

    Download ADCP Data (os75 netcdf) in2019_e01_os75nb_nc.zip

    Download overview plots of collected data: in2019_e01_os75nb_webpy.zip

    Download ADCP Data (os150 netcdf) in2019_e01_os150nb_nc.zip

    Download overview plots of collected data: in2019_e01_os150nb_webpy.zip

    Multibeam echosounder
    Time (UTC): to
    Latitude: 0.00
    Longitude: 0.00
    Datafiles: 215
    Total size: 53.01 Gb
    Get Data from Trawler

    Radon measurements
    Time (UTC): 2019-08-01 00:00 to 2019-08-02 23:00
    Latitude: -16.93 to -13.89
    Longitude: 145.78 to 146.27
    Measurements:36
    Get Data from Trawler

    Download RV Investigator Radon dataset (XLSX)

    Sub-bottom profiler
    Time (UTC): 2019-07-27 13:07 to 2019-08-02 23:08
    Latitude: -16.84 to -9.22
    Longitude: 133.24 to 146.34
    Get Data from Trawler

    Geophysical Survey and Mapping data acquisition and processing report

    Singlebeam echosounder
    Time (UTC): 2019-07-27 13:07 to 2019-08-02 23:08
    Latitude: -16.84 to -9.22
    Longitude: 133.24 to 146.34
    Get Data from Trawler

    Geophysical Survey and Mapping acquisition report

    XBT Profile
    Time (UTC): 2019-07-27 11:07 to 2019-07-30 09:07
    Latitude: -9.94 to -9.23
    Longitude: 133.13 to 144.83
    Quantity: 4 Casts
    Max Pressure: 1100.3 db
    Get Data from Trawler

    Caveat: this event log was authored on the voyage using either a standalone logging system called ELOG or more recently an inhouse event logging tool called Everlog. For this reason, some positions are unknown and may be resolved using the logged time and the underway dataset for the voyage.

    Search for events from this voyage and others

    Events ordered by start time:- Click on column header to sort. Use link on 'When' to see full details. Use the CSV button at the bottom to export data.

    When (UTC) LatitudeLongitude InstrumentIdentifierActionLogSubject / Comments
    2019-07-21 09:08 [view ] SMPS [view] Maintenance in2019_e01_Atmospherics Running overnight zero, starting at measurement 15 or 10:17am
    2019-07-21 10:18 [view ] CCNC [view] Maintenance in2019_e01_Atmospherics CCN calibration Set-up CCN calibration chassis. Power on CCNC, dry start up mode. All OK.
    2019-07-21 15:00 [view ] CPC 3772 [view] Maintenance in2019_e01_Atmospherics Annual maintenance Maintenance on the CPC including: - Butanol fill filter replacement, including regeneration of the old filter after ultrasonic bath - Nozzle and OPC clean. Nozzle was pretty clean. OPC had a bunch of white stuff in it which was cleaned out with a Q-tip and water. Noted there was a bit of uneven grease on the o-ring between the column and OPC. - Inspected internal plumbing - all connections were good, and no butanol was in the wrong lines. - Wick was removed and dried overnight in the fume cupboard, reinstalled the next morning all good.
    2019-07-21 15:00 [view ] SMPS [view] Maintenance in2019_e01_Atmospherics Annual maintenance SMPS maintenance including: - Cleaning the external impactors with kimwipe and alcohol - the nozzle on the impactor (not the preimpactor) may have been installed backwards. See new procedure for proper installation. - Cleaning the impactor at the top of the column with kimwipe and alcohol. - Reapplicaiton of a thin layer of grease to all the impactors - Cleaning of the DMA and all surfaces with kimwipes and alcohol. The DMA itself was very clean, and it was just the sections around the external screws that had a bit of corrosion on them. The brass connections looked reasonable, however one side (presumably the inlet) was dirty on the inside. - Changed the excess air filter with a new one.
    2019-07-22 01:41 [view ] SMPS [view] Maintenance in2019_e01_Atmospherics Overnight zero check Completed overnight zero check. Total concentrations averaged about 5e-1, with a max of 1.5e0. This is acceptable.
    2019-07-22 02:18 [view ] SMPS [view] Maintenance in2019_e01_Atmospherics Flow check Gilibrator flow check - 293.1 ccm from 6 points. Only small deviation around this of approx +/- 1 ccm. Goal is 300 ccm, so this is acceptable within 5%.
    2019-07-22 02:20 [view ] SMPS [view] Maintenance in2019_e01_Atmospherics SMPS operation Leaving SMPS on, measuring lab air, until the PSL cal is performed in a few days.
    2019-07-22 08:59 [view ] CPC 3772 [view] Maintenance in2019_e01_Atmospherics Zero check started Re-installed instrument after 12M service, positioned next to CCN inlet. Calibrated flows (gilibrator = 999ccm (n=5), Alicat set= 1.000LPM, Alicat read= 1.000LPM) HEPA zero check started 0900hrs UTC.
    2019-07-23 01:17 [view ] CPC 3772 [view] Stop in2019_e01_Atmospherics Zero check stopped overnight zero check stopped 0120hrs UTC. CPC back measuring lab air.
    2019-07-23 01:50 [view ] Nephelometer [view] Maintenance in2019_e01_Atmospherics Nephelometer 12 month maintenance including: - Exhaust ball valve cleaning - inlet ball valve was clean, but exhaust ball valve was quite dirty. - Optic cell clean performed - this was quite clean. - Full optical bench clean. A small amount of vac grease was applied to o-rings in optic chamber tubes. - Inspected the exhaust flow DFU - all OK, none replaced. - Batteries were replaced. - Pneumatics were cleaned After reassembly, a leak check was performed - Failed Isolated the external plumbing from the leak test - still failed. Discovered tubing to the pressure sensor was pinched during optic bench reinstallation - test still failed. Found a tube completely disconnected on the output of the zero pump, which means the zero tests for the foreseeable past are crap, but will need to look into the past zeros to verify data quality. This is an inherent problem with the design of these connections.
    2019-07-27 09:12 [view ] CCNC [view] Maintenance in2019_e01_Atmospherics supersaturation settting I had the CCNc in & 39;manual& 39; SS mode for a few hrs this afternoon. back on auto ss 09:15am
    2019-07-27 09:28 [view ] -9.1165 132.7817 Techsas1 [view] Start in2019_e01 started acquisition upon entering australian EEZ
    2019-07-27 09:28 [view ] -9.1165 132.7817 Techsas2 [view] Start in2019_e01 started acquisition upon entering australian EEZ
    2019-07-27 09:53 [view ] -9.1415 132.8613 EM710 [view] Start GSM_in2019_e01 Start pinging
    2019-07-27 09:54 [view ] -9.1423 132.8638 EM122 [view] Start GSM_in2019_e01 Start pinging
    2019-07-27 10:23 [view ] -9.1725 132.9593 XBT [view] Crash GSM_in2019_e01 XBT failure- not connecting to hardware- bad connection wire
    2019-07-27 11:16 [view ] -9.2253 133.1262 XBT [view] Drop GSM_in2019_e01 XBT drop after fixed by using splitter at launch connection SST=26.204
    2019-07-27 11:32 [view ] -9.2415 133.1778 EM710 [view] SVP GSM_in2019_e01 New SVP applied in2019_e01_xbt_001.asvp- start logging
    2019-07-27 11:33 [view ] -9.2420 133.1792 EM122 [view] SVP GSM_in2019_e01 New SVP applied in2019_e01_xbt_001.asvp- start logging
    2019-07-27 11:35 [view ] -9.2445 133.1872 EK60 [view] Start GSM_in2019_e01 Start pinging/logging all frequencies to 1500m using EK80 software
    2019-07-27 11:36 [view ] -9.2450 133.1887 SBP [view] Start GSM_in2019_e01 Start logging
    2019-07-27 11:50 [view ] -9.2597 133.2352 EK60 [view] Start GSM_in2019_e01 All frequencies active- previously recording in passive
    2019-07-27 12:28 [view ] -9.2977 133.3560 EM710 [view] GSM_in2019_e01 Testing applying Seapath with same offsets to check for motion artefact
    2019-07-27 14:00 [view ] -9.3913 133.6527 Gravity Meter [view] Start GSM_in2019_e01 Gravity meter started Air shocks all filled to 30 psi
    2019-07-27 14:47 [view ] -9.4390 133.8038 Gravity Meter [view] Stop GSM_in2019_e01 Gravity meter shut back down- drifting after unclamping and ST tracking on
    2019-07-27 15:28 [view ] -9.4810 133.9365 Gravity Meter [view] Start GSM_in2019_e01 Trying to start gravity meter again
    2019-07-27 21:42 [view ] -9.8583 135.1333 ADCP [view] in2019_e01 Stop logging for multibeam noise test while slowing for seismic deployment
    2019-07-27 21:43 [view ] -9.8592 135.1358 EM710 [view] Stop GSM_in2019_e01 All acoustics stopped for noise testing em's
    2019-07-27 22:18 [view ] -9.8755 135.1897 EM710 [view] GSM_in2019_e01 All acoustics back on following noise tests
    2019-07-27 22:20 [view ] -9.8760 135.1908 ADCP [view] in2019_e01 Start
    2019-07-27 23:34 [view ] -9.8958 135.2510 SMPS [view] Maintenance in2019_e01_Atmospherics Flow check Gilibrator flow check at inlet - 0.280 LPM
    2019-07-27 23:37 [view ] -9.8970 135.2548 Underway Seawater Data Start in2019_e01 Underway Data Collection Started - optode 662 installed coefficients updated IAW .xls file
    2019-07-27 23:39 [view ] -9.8977 135.2572 Other [view] Start GSM_in2019_e01 Start logging qinsy for seismic trials
    2019-07-28 00:16 [view ] -9.9117 135.3018 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_002.asvp applied
    2019-07-28 00:36 [view ] -9.9207 135.3300 SMPS [view] Maintenance in2019_e01_Atmospherics butanol filled and emptied
    2019-07-28 01:50 [view ] -9.9438 135.4037 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_003.asvp applied
    2019-07-28 02:16 [view ] -9.9538 135.4350 SMPS [view] Fault in2019_e01_Atmospherics Condenser temperature not reaching setpoint Butanol ran out (error 40) overnight. I refiled it and the signal came back good but error 32 is persisting (temperatures not reaching setpoint). I tried restarting the software and running the warmup, but the condenser temperature is not reaching the setpoint (10oC), and only making it to 14oC. I power cycled the instrument, leaving it off for about 30 minutes, then unplugging it from mains power for about 1 minute, before restarting (without the software). Checked about 45 min later, and the condenser temperature is down to 11.8oC. Leaving it for a while longer to see if it gets down to temperature. Left another 45 min, temperature hasn& 39;t gone any further down. Took cover off to inspect plumbing, nothing suspicious, but the air conditioner is blowing directly on the instrument - maybe the heater dump for the temperature isn& 39;t able to work against the 21oC AC. I& 39;ve readjusted the AC flow direction flaps to point directly out, rather than down as they were. No change after 15 minutes. It could be related to the humidity or some dodgy butanol - I& 39;ve connected the diffusion drier to the inlet (still sampling lab air) to test the humidity effect. No change. Leaving overnight to see if the wick wetting is the issue (removed the diffusion drier).
    2019-07-28 02:24 [view ] -9.9570 135.4448 MAAP [view] Maintenance in2019_e01_Atmospherics Annual maintenance Opened up optics to discover the reflection optics are very soiled (see attached image) Need to fully disassemble optics to clean out properly in an IPA bath. The transmission optics have been discovered to be quite dirty as well.(see attached image) Removal of the stage has discovered that silicon has been applied in a poor manner to keep a seal on, but this has overflowed to stick the stage on in the back left corner. Cleaned up and removed this excess silicon. Removed the "O-ring" seal, and brass shim surrounding the transmission optics. Can& 39;t remove the transmission optics itself, so we& 39;re cleaning it in place with canned IPA and Q-tips making sure to clean the diode surfaces too. Cleaning the reflection optics in a bath of IPC and agitating Disassembling all the transmission optics and fully cleaning in ultrasonic bath, lots of rust everywhere. After cleaning, discovered some of the matte black paint on the stage has been removed - this puts into question the quality of the data going forward. Need to check data going forward. After full cleaning in bath, and rust removal with glass fibre pen and lots of q-tips, everything looked good. But after a while in the oven, Ian discovered liquid still in the reflection optics head. We took it apart a bit more to remove the excess water and while apart, put it in the oven for a few more hours.
    2019-07-28 02:48 [view ] -9.9658 135.4732 EK60 [view] Other GSM_in2019_e01 Calibration values added to ek80 software
    2019-07-28 03:57 [view ] -10.0010 135.5862 Air Sampling Pump Maintenance in2019_e01_Atmospherics would not start - cleaned Pump would not start. Inspection found it filthy inside. I cleaned it really thoroughly, reassembled, and now it works
    2019-07-28 03:59 [view ] -10.0023 135.5903 Air Sampling Pump Start in2019_e01_Atmospherics started - frequency adjusted Flow rate measured 350LPM on restarted following maintenance. Normal flow rate 420LPM frequency adjusted (from 30Hz, to 36Hz). Flow now 420LPM
    2019-07-28 04:01 [view ] -10.0047 135.5977 Air Sampling Inlet Flowmeter Start in2019_e01_Atmospherics Started following dry dock
    2019-07-28 04:03 [view ] -10.0067 135.6043 Air Sampling Inlet Flowmeter Start in2019_e01_Atmospherics Started following dry dock
    2019-07-28 04:57 [view ] -10.0657 135.7915 CCNC [view] Change in2019_e01_Atmospherics CCNc supersaturation SS set to manual mode 04:57UTC hrs
    2019-07-28 05:16 [view ] -10.0858 135.8555 EM122 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_004.asvp applied
    2019-07-28 05:17 [view ] -10.0863 135.8572 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_004.asvp applied
    2019-07-28 05:52 [view ] -10.1250 135.9800 Nephelometer [view] Maintenance in2019_e01_Atmospherics Running overnight zero, starting at 05:50am UTC
    2019-07-28 06:38 [view ] -10.1743 136.1368 CCNC [view] Change in2019_e01_Atmospherics supersaturation settting changed back to auto mode 06:30hrs
    2019-07-28 06:42 [view ] -10.1790 136.1513 SMPS [view] Fault in2019_e01_Atmospherics Power cycled the instrument to try to get condenser temp to set point This helped it get down to 11.2oC now after 45 minutes. Power cycling again to see if I can get it that little bit further!
    2019-07-28 21:02 [view ] -11.0430 138.8995 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_005.asvp applied
    2019-07-28 21:51 [view ] -11.0927 139.0580 CCNC [view] Change in2019_e01_Atmospherics Zero check and SS HEPA zero check 09:45- 09:50pm switched to manual SS 09:53hrs for todays SS calibration.
    2019-07-29 00:01 [view ] -11.1717 139.4788 SMPS [view] Fault in2019_e01_Atmospherics Condenser temperature still out of range Came in in the morning to find the condenser temperature at 15 degrees. The software has been having trouble connecting, so maybe this is a standby mode feature? I have restarted the instrument and the PC to see if I have any luck.
    2019-07-29 00:30 [view ] -11.1552 139.5783 pCO2 [view] error in2019_e01 Reference gas cylinders were not open until now An oversight on my behalf, the gas bottles were not opened from when the underway system was started until now.
    2019-07-29 00:49 [view ] -11.1443 139.6438 Underway Seawater Data Maintenance in2019_e01 Dosing Pump install was tested with the underway line & 39;dosed& 39; with approximately 10L of 10% Fresh water over apprximately 10 minutes.
    2019-07-29 01:27 [view ] -11.1222 139.7778 SMPS [view] Fault in2019_e01_Atmospherics Condenser temp I& 39;ve pulled the SMPS out in order to dry the wick. Disconnected the DMA to pull it out. Taken it to dry lab. Disconnected the CPC (lots of cables!) and removed it (4 screws underneath) and relocated the CPC to the fume hood. Removed the top cover of the optics, one of the tiny circuit boards blocking one of the screws for the front plate, then removed the front plate and foam, removed the inlet protecting the saturator block and then removed the wick with tweezers. The wick is a slightly yellow colour, not sure if it was originally white or what, but it may be discoloured. Leaving in the fume hood for the day. Might put in the oven overnight to drive the last bit of the water out.
    2019-07-29 05:11 [view ] -10.9952 140.5470 CCNC [view] Change in2019_e01_Atmospherics supersaturation settting changed SS to deltaT = 14.0 (SS=0.97) at 05:10hrs
    2019-07-29 05:27 [view ] -10.9867 140.5987 MAAP [view] Observation in2019_e01_Atmospherics filter tape filter tape damaged on left hand side (new side)? swiching instrument off until someone can fix it.
    2019-07-29 05:51 [view ] -10.9728 140.6818 MAAP [view] Maintenance in2019_e01_Atmospherics Fixed filter tape damage Stopped MAAP, opened up the head, cut out the damaged filter tape and taped the two parts together and wrapped it around a few times. Looks good now. Currently zeroing.
    2019-07-29 08:42 [view ] -10.8602 141.2643 Picarro Spectrometer [view] Start in2019_e01_Atmospherics Picarro started, calibrated, running monitor runs I have reinstalled the Picarro following dry dock. I have run calibrations for over 3 hours, and now runnning monitor runs
    2019-07-29 10:18 [view ] -10.7607 141.5847 MAAP [view] Outage in2019_e01_Atmospherics Restarted to clear error after switching from manual mode.
    2019-07-29 10:20 [view ] -10.7585 141.5920 Nephelometer [view] Maintenance in2019_e01_Atmospherics Zero check stopped, started back on ambient
    2019-07-29 10:21 [view ] -10.7577 141.5947 Change in2019_e01_Atmospherics CPC and SMPS taps were open for the past few days. Closed them now, so no longer sampling lab air.
    2019-07-29 10:23 [view ] -10.7560 141.5998 CCNC [view] Maintenance in2019_e01_Atmospherics CCN cal finished, set back to auto sampling, connected back up to main inlet CCN tap was also open, so this is now the start of proper ambient sampling on that line.
    2019-07-29 10:32 [view ] -10.7460 141.6315 Fluorometer [view] Cleaned in2019_e01
    2019-07-29 20:16 [view ] -9.7960 143.1838 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_006.asvp applied
    2019-07-29 23:57 [view ] -9.3263 143.7108 CCNC [view] Change in2019_e01_Atmospherics supersaturation settting SS setting control was still in & 39;manual& 39; mode, after yesterdays calibration (SS= 0.09%, Dt= 3.0). Switched to Auto SS 12:00am UTC
    2019-07-30 00:02 [view ] -9.3165 143.7260 CPC 3772 [view] Observation in2019_e01_Atmospherics 3772 Alicat MFC Alicat remote screen off/not working? Any keypad press does not activate it. So cannot confirm flows. Is it going into a sleep mode?
    2019-07-30 01:02 [view ] -9.2247 143.8993 Nephelometer [view] Maintenance in2019_e01_Atmospherics Added missing insulation to inlet pipe that was missed during servicing.
    2019-07-30 01:04 [view ] -9.2252 143.9088 pCO2 [view] in2019_e01 Adjusted flow rate from STD4
    2019-07-30 01:38 [view ] -9.2322 144.0270 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_007.asvp applied
    2019-07-30 03:22 [view ] -9.3438 144.3692 Aerosol Manifold [view] Maintenance in2019_e01_Atmospherics Air port open for last 20mins
    2019-07-30 03:43 [view ] -9.4027 144.4167 MAAP [view] Change in2019_e01_Atmospherics BC concentration unit Changed displayed [BC] units from ug/m3 to ng/m3 at 03:45am UTC. The displayed concentration on the front screen was constantly 0.00 ug/m3. Post-change [BC] = 17ng/m3.
    2019-07-30 04:06 [view ] -9.4642 144.4663 SMPS [view] Maintenance in2019_e01_Atmospherics PSL cal at 303 +/- 6 nm started Condenser temperature at 18.2oC (set point of 10oC).
    2019-07-30 04:14 [view ] -9.4860 144.4840 SMPS [view] Maintenance in2019_e01_Atmospherics PSL cal at 100 +/- 3 nm started
    2019-07-30 04:17 [view ] -9.4860 144.4840 SMPS [view] Maintenance in2019_e01_Atmospherics PSL cal at 508 +/- 8 nm started
    2019-07-30 04:21 [view ] -9.4860 144.4840 SMPS [view] Maintenance in2019_e01_Atmospherics PSL cal at 81 +/- 3 nm started
    2019-07-30 04:21 [view ] -9.5032 144.4978 Nephelometer [view] Maintenance in2019_e01_Atmospherics Temperature, BaroPress, RH sensor calibration Audited sample temperature, Barometric pressure, and sample RH sensors against external probes (about 2hrs ago). All passsed, no adjustment performed. Parameter Nephelometer Calibrated reference Temperature 25.69 26.34 Rel Humidity 44.775 44.93 Pressure 1010.1 1009
    2019-07-30 04:24 [view ] -9.5110 144.5042 Nephelometer [view] Maintenance in2019_e01_Atmospherics leak check performed instrument leak check, from software settings (exhaust plumbed to inlet). all passed.
    2019-07-30 04:30 [view ] -9.5265 144.5167 Nephelometer [view] Start in2019_e01_Atmospherics Zero check
    2019-07-30 04:52 [view ] -9.5828 144.5622 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_08.asvp applied
    2019-07-30 05:12 [view ] -9.6367 144.6057 Nephelometer [view] Start in2019_e01_Atmospherics Span check
    2019-07-30 05:19 [view ] -9.6543 144.6200 pCO2 [view] Filter Clean in2019_e01
    2019-07-30 05:20 [view ] -9.6552 144.6207 Fluorometer [view] Cleaned in2019_e01
    2019-07-30 05:20 [view ] -9.6557 144.6210 TSG [view] Filter Clean in2019_e01
    2019-07-30 05:45 [view ] -9.7085 144.6832 EM122 [view] Start GSM_in2019_e01 Start logging
    2019-07-30 05:45 [view ] -9.7090 144.6842 EM710 [view] GSM_in2019_e01 SIS grid not working
    2019-07-30 05:54 [view ] -9.7237 144.7098 EM122 [view] GSM_in2019_e01 SIS offsets zeroed
    2019-07-30 05:55 [view ] -9.7248 144.7118 EM710 [view] GSM_in2019_e01 SIS offsets zeroed, start new survey D
    2019-07-30 05:59 [view ] -9.7312 144.7228 pCO2 [view] Adjusted in2019_e01 Flow rate from STD4 is now correct. STD 1-3 decreased IAW email from John Akl
    2019-07-30 06:11 [view ] -9.7532 144.7517 XBT [view] Drop GSM_in2019_e01 XBT drop 2 prior to EM calibration site
    2019-07-30 06:18 [view ] -9.7598 144.7603 Nephelometer [view] Maintenance in2019_e01_Atmospherics Full calibration activated
    2019-07-30 06:19 [view ] -9.7600 144.7607 XBT [view] Drop GSM_in2019_e01 XBT drop 3
    2019-07-30 06:21 [view ] -9.7608 144.7628 MAAP [view] Change in2019_e01_Atmospherics BC concentration unit after speaking to Ian McRobert, changed [BC] back to ug/m3 until all data processes downstream are verified and notified.
    2019-07-30 06:36 [view ] -9.7708 144.7733 Other [view] SVP GSM_in2019_e01 Rapid cast probe deployed from net drum
    2019-07-30 06:37 [view ] -9.7710 144.7737 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_xbt_003.asvp applied
    2019-07-30 06:37 [view ] -9.7710 144.7737 EM122 [view] SVP GSM_in2019_e01 New SVP in2019_e01_xbt_003.asvp applied
    2019-07-30 06:46 [view ] -9.7725 144.7743 EM710 [view] Other GSM_in2019_e01 Frequency range changed to 40-100kHz Initial issue changing- faulty ping status and owuld not transfer settings. Stopped and started pinging again, eventually came in. Issues tracking bottom in stbd sector
    2019-07-30 06:48 [view ] -9.7735 144.7748 SBP [view] Stop GSM_in2019_e01 Stop pinging for EM cal
    2019-07-30 07:16 [view ] -9.7688 144.7485 EM710 [view] Start GSM_in2019_e01 Roll line 1 Max angles 66/66 40-100kHz Dynamic
    2019-07-30 07:17 [view ] -9.7708 144.7485 EM122 [view] Start GSM_in2019_e01 Roll line 1 Max angles 66/66
    2019-07-30 07:20 [view ] -9.7760 144.7485 EM710 [view] Start GSM_in2019_e01 Pitch line 1 Beam angles 40/40
    2019-07-30 07:20 [view ] -9.7772 144.7485 EM122 [view] Start GSM_in2019_e01 Pitch line 1 Beam angles 40/40
    2019-07-30 07:34 [view ] -9.7907 144.7482 Nephelometer [view] Maintenance in2019_e01_Atmospherics Span check post-adjust Span check
    2019-07-30 07:41 [view ] -9.7762 144.7485 EM710 [view] Start GSM_in2019_e01 Roll line 2 Max angles 66/66
    2019-07-30 07:41 [view ] -9.7762 144.7485 EM122 [view] Start GSM_in2019_e01 Roll line 2 Max angles 66/66
    2019-07-30 07:54 [view ] -9.7722 144.7545 EM710 [view] Start GSM_in2019_e01 Start yaw line 1 Angles 15/72
    2019-07-30 07:54 [view ] -9.7722 144.7545 EM710 [view] Start GSM_in2019_e01 Start yaw line 1 Angles 15/68
    2019-07-30 08:11 [view ] -9.7943 144.7415 Nephelometer [view] Maintenance in2019_e01_Atmospherics Zero check Post-adjust Zero Check
    2019-07-30 08:22 [view ] -9.7737 144.7423 EM710 [view] Start GSM_in2019_e01 Yaw line 2 Beam angles 15/72
    2019-07-30 08:22 [view ] -9.7727 144.7423 EM122 [view] Start GSM_in2019_e01 Yaw line 2 Beam angles 15/68
    2019-07-30 08:32 [view ] -9.7763 144.7483 EM710 [view] Start GSM_in2019_e01 Start latency line Beam angles 40/40 Spped 3 kts
    2019-07-30 08:32 [view ] -9.7763 144.7483 EM122 [view] Start GSM_in2019_e01 Start latency line Beam angles 40/40 Spped 3 kts
    2019-07-30 08:44 [view ] -9.7860 144.7487 Nephelometer [view] Observation in2019_e01_Atmospherics Zero check Zero cal/check does not appear to be working proplery. Cal ball valves did not close, cal control box did not work, zero air pump did not switch on. Will investigate tomorrow.
    2019-07-30 09:24 [view ] -9.8572 144.7975 Nephelometer [view] Maintenance in2019_e01_Atmospherics Zero Check test tested zero check functionality via remote terminal /udpterminal
    2019-07-30 10:04 [view ] -9.9720 144.8257 XBT [view] Drop GSM_in2019_e01 XBT drop 4 prior to deep water cal site
    2019-07-30 10:09 [view ] -9.9860 144.8242 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_xbt_004.asvp applied
    2019-07-30 10:09 [view ] -9.9860 144.8242 EM122 [view] SVP GSM_in2019_e01 New SVP in2019_e01_xbt_004.asvp applied
    2019-07-30 10:33 [view ] -9.9933 144.8513 EM710 [view] Start GSM_in2019_e01 Start deep cal roll 1 Beam angles 66/66 (reality getting 47/44)
    2019-07-30 10:33 [view ] -9.9933 144.8513 EM122 [view] Start GSM_in2019_e01 Start deep cal roll 1 Beam angles 66/66
    2019-07-30 10:38 [view ] -9.9905 144.8418 Underway Steve Van Graas Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-30 10:39 [view ] -9.9898 144.8395 EM710 [view] Start GSM_in2019_e01 Pitch line 1 Beam angles 40/40
    2019-07-30 10:39 [view ] -9.9898 144.8395 EM122 [view] Start GSM_in2019_e01 Pitch line 1 Beam angles 40/40
    2019-07-30 10:42 [view ] -9.9880 144.8340 Underway steve vg Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-30 10:44 [view ] -9.9868 144.8307 Underway Stephanie Zeliadt Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-30 11:02 [view ] -9.9810 144.8135 EM710 [view] Start GSM_in2019_e01 Start pitch line 2
    2019-07-30 11:02 [view ] -9.9810 144.8135 EM122 [view] Start GSM_in2019_e01 Start pitch line 2
    2019-07-30 11:31 [view ] -9.9975 144.8350 EM710 [view] Start GSM_in2019_e01 Start yaw line 1 Max angles 15/72
    2019-07-30 11:31 [view ] -9.9975 144.8350 EM122 [view] Start GSM_in2019_e01 Start yaw line 1 Max angles 15/72 Heading 288
    2019-07-30 11:53 [view ] -9.9725 144.8162 EM710 [view] Start GSM_in2019_e01 Deep yaw line 2
    2019-07-30 11:53 [view ] -9.9725 144.8162 EM122 [view] Start GSM_in2019_e01 Deep yaw line 2- offline Heading 108
    2019-07-30 12:12 [view ] -9.9795 144.8367 EM710 [view] Start GSM_in2019_e01 Start deep yaw 3 heading 288
    2019-07-30 12:12 [view ] -9.9795 144.8367 EM122 [view] Start GSM_in2019_e01 Start deep yaw 3 heading 288 Max angles 72/15
    2019-07-30 12:22 [view ] -9.9730 144.8173 EM710 [view] GSM_in2019_e01 End of patch test- all angles back to 72
    2019-07-30 14:09 [view ] -10.2810 144.9432 EK60 [view] Start GSM_in2019_e01 Start pinging/logging using ER60 software
    2019-07-30 14:19 [view ] -10.3127 144.9528 SBP [view] Start GSM_in2019_e01 Start pinging/logging
    2019-07-30 14:28 [view ] -10.3395 144.9612 EM710 [view] Stop GSM_in2019_e01 Stop logging
    2019-07-30 16:05 [view ] -10.6415 145.0545 Underway richard Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-30 16:07 [view ] -10.6503 145.0572 Underway Richard Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-30 16:09 [view ] -10.6548 145.0587 Underway steve vg Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-30 16:15 [view ] -10.6745 145.0647 ADCP [view] Start in2019_e01 Start Logging following patch test completion
    2019-07-31 00:36 [view ] -11.9413 145.4645 MAAP [view] Maintenance in2019_e01_Atmospherics Disconnected MAAP sample line for maintenance MAAP disconnected 12:20 hrs UTC
    2019-07-31 00:38 [view ] -11.9412 145.4645 MAAP [view] Maintenance in2019_e01_Atmospherics flow audit & 39;as is& 39; flow audit, gilibrator drawing lab air, T1 and Tref co-located Gilibrator = 16.52 (n=5), MAAP = 16.67; error = 0.6% T1 = 24.8, Tref -= 23.8 MAAP baroP = 1014, Ref BaroP = 1014 ---------------------------------------------------------------- Adjustments: T1 adjusted - 1 BaroP unchanged -------------------------------------------------------------- Final checks: Gilibrator = 16.62 (n=5), MAAP = 16.67 LPM; error 0.3% T1 = 23.3, Tref = 23.4
    2019-07-31 02:36 [view ] -11.9412 145.4645 MAAP [view] Observation in2019_e01_Atmospherics MAAP Ozone interference test stopped MAAP disconnected from Ozone calibrator and reconnected to aerosol sample line 02:35hrs. No evidence of ozone interference found.
    2019-07-31 02:38 [view ] -11.9413 145.4643 Aerosol Manifold [view] Change in2019_e01_Atmospherics MAAP ball valve open Aerosol manifold possibly drawing in air from Aerosol Lab ~01:00 to 02:40hrs. MAAP had been disconnected from it& 39;s nafion dryer for testing, but MAAP inlet ball valve had not been closed. So possibly some back-flow into aerosol manifold during this time.
    2019-07-31 03:24 [view ] -11.9405 145.4648 Nephelometer [view] Maintenance in2019_e01_Atmospherics flow audit Neph disconnected from sample line 02:30 - 03:20hrs for flow calibration. Flows measured at inlet of Neph. Gilib (initial) = 5.121LPM (n=5). Alicat set = 5.000slpm Gilib (final) = 5.019LPM (n=5). Alicat set= 4.950slpm, read= 4.950slpm/14.76(v)lpm.
    2019-07-31 04:37 [view ] -11.9408 145.4645 CTD [view] 1 in2019_e01 CTD Frame Only Deployed for Ship Winch Test
    2019-07-31 10:05 [view ] -11.9570 145.4817 ADCP [view] in2019_e01 OS150 and OS75 turned off for multibeam noise testing
    2019-07-31 10:07 [view ] -11.9567 145.4817 EM122 [view] Start GSM_in2019_e01 Start noise test
    2019-07-31 10:12 [view ] -11.9552 145.4807 Underway Stephanie Zeliadt Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-31 10:13 [view ] -11.9548 145.4805 Underway steve vg Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-31 10:16 [view ] -11.9533 145.4805 Underway Stephanie Zeliadt Hydrochem Lab : Sample TSG_Samples_in2019_e01
    2019-07-31 11:19 [view ] -12.0610 145.5002 EM122 [view] Stop GSM_in2019_e01 End noise tests
    2019-07-31 11:19 [view ] -12.0625 145.5005 EM122 [view] Start GSM_in2019_e01 Start pinging/logging New Patch values applied from this point forward
    2019-07-31 14:07 [view ] -12.5253 145.6390 EM122 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_09.asvp applied
    2019-07-31 22:19 [view ] -13.8878 146.0653 CTD [view] 001 Deployed in2019_e01
    2019-07-31 22:32 [view ] -13.8880 146.0650 CTD [view] Retrieved in2019_e01
    2019-07-31 23:49 [view ] -13.8880 146.0650 Radon Detector [view] Start in2019_e01_Atmospherics Started following dry dock
    2019-07-31 23:50 [view ] -13.8880 146.0650 Radon Detector [view] Maintenance in2019_e01_Atmospherics Calibration started Calibration started. Flush for 6 hours then cal for 6 hours. Note we do not normally flush, as the dry air is constantly flushiing the source. The flush was added in this instance as an extra precaution
    2019-08-01 00:26 [view ] -13.8877 146.0652 SMPS [view] Outage in2019_e01_Atmospherics Switched instrument off to packup for repair Error 32 suggests that the peltier cooler is broken and GRIMM are asking to send it back to them in Germany. Ian is going to try to fix it in Hobart before sending it back, so I& 39;m packing everything up, including the PC.
    2019-08-01 01:34 [view ] -13.8872 146.0658 CPC 3772 [view] Maintenance in2019_e01_Atmospherics Flow cal Set the CPC back up and reconfigured the plumbing so that the orifice pressure on the CPC will register OK. Connected up to the strong vacuum so that the orifice pressure reads OK (this is fine because the MFC is maintaining the flow actual flow). When doing the flow cal with lab air: gilibrator 1.001 LPM set point 0.990 LPM (volumetric) When doing the flow cal inline: gilibrator 1.005 LPM set point 1.000 LPM The inline set point is really variable, but after an average of 10 tests, we& 39;ve decided to leave it as above.
    2019-08-01 01:41 [view ] -13.8872 146.0658 CPC 3772 [view] Start in2019_e01_Atmospherics CPC start, reconnected to main line
    2019-08-01 03:16 [view ] -13.8882 146.0643 Both Stop in2019_e01_VSAT Stopped for mast maintenance
    2019-08-01 03:17 [view ] -13.8882 146.0643 Both Start in2019_e01_VSAT Resumed normal operation post maintenance
    2019-08-01 03:17 [view ] -13.8882 146.0645 Both Auto Trim in2019_e01_VSAT Auto trim after vsat restart
    2019-08-01 03:36 [view ] -13.8878 146.0647 Fluorometer [view] Cleaned in2019_e01
    2019-08-01 04:31 [view ] -13.8895 146.0642 Other [view] GSM_in2019_e01 Drop keels flush with hull
    2019-08-01 04:31 [view ] -13.8897 146.0642 Other [view] Start GSM_in2019_e01 POSMV GAMS calibration- starting figure 8s
    2019-08-01 06:00 [view ] -14.0207 146.0998 Other [view] GSM_in2019_e01 End GAMS
    2019-08-01 06:46 [view ] -14.1567 146.1433 EM122 [view] Start GSM_in2019_e01
    2019-08-01 09:19 [view ] -14.6442 146.2222 EK60 [view] Start GSM_in2019_e01 Start pinging/logging to 1500m with EK80 software
    2019-08-01 11:43 [view ] -15.0993 146.2958 Other [view] in2019_e01 restarted DC1 VM
    2019-08-01 14:16 [view ] -15.6043 146.3285 EM122 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_010.asvp applied
    2019-08-01 18:28 [view ] -16.4490 146.2613 ADCP [view] Start in2019_e01
    2019-08-01 20:18 [view ] -16.7440 146.1480 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_sst_011.asvp applied
    2019-08-01 21:37 [view ] -16.8172 145.9818 ADCP [view] Stop in2019_e01 Stopped for reference surface work GSM
    2019-08-01 21:39 [view ] -16.8150 145.9800 Other [view] Stop GSM_in2019_e01 Stop all systems except em710 for reference surface
    2019-08-01 22:23 [view ] -16.7652 145.9403 EM710 [view] SVP GSM_in2019_e01 New SVP in2019_e01_svp_001.asvp applied
    2019-08-01 22:29 [view ] -16.7695 145.9445 EM710 [view] Start GSM_in2019_e01 AHO Reference Surface Line 1
    2019-08-01 22:43 [view ] -16.8725 145.8078 Underway Seawater Data Stop in2019_e01 Underway Pumps Shut Down for End of E01
    2019-08-01 22:44 [view ] -16.7805 145.9520 EM710 [view] Start GSM_in2019_e01 AHO Reference Surface Line 2
    2019-08-01 23:02 [view ] -16.7743 145.9498 EM710 [view] Start GSM_in2019_e01 AHO Reference Surface XLine 1 Slightly crabbing at start
    2019-08-01 23:21 [view ] -16.7975 145.9238 ADCP [view] Start in2019_e01
    2019-08-02 00:08 [view ] -16.8132 145.8402 CCNC [view] Maintenance in2019_e01_Atmospherics Empty drain, fill water supply, zero check.
    2019-08-02 00:12 [view ] -16.8240 145.8360 CCNC [view] Start in2019_e01_Atmospherics Resumed ambient sampling
    2019-08-02 00:13 [view ] -16.8267 145.8347 CPC 3772 [view] Maintenance in2019_e01_Atmospherics Emptied butanol catchpot, needed to unplug the vacuum to allow to drain.
    2019-08-02 01:58 [view ] -16.9257 145.7807 Voyage Stop in2019_e01 Stopped techsas and finalised voyage.

    Save this list as

    Deployments by start time:- Click on column header to sort. Use link on 'Deployment' to see full details. Use the CSV button at the bottom to export data (tab delimited).

    Search for deployments from this voyage and others

    Deployment OperationStationCastGearProject Start and end timesStart LocationEnd LocationDepthDescription
    ADCP Data [details]   Teledyne RDI ADCP 150 kHz,Teledyne RDI ADCP 75 kHz 2457 2019-07-27 20:25 2019-08-02 11:56 -9.1745 132.9656 -16.9258 145.7807 782
    XBT Cast [details]   2391 2019-07-27 21:16 404.9
    XBT Cast [details]   2391 2019-07-30 16:11 1100.3
    XBT Cast [details]   2391 2019-07-30 16:19 1100.3
    XBT Cast [details]   2391 2019-07-30 19:55 996.8

    Save this list as

    NCMI Information and Data Centre  »  Applications  »  Data Trawler